OpenStudy (anonymous):

what is the laplace transform of y'' - y' -2y = -8cost-2sint, with initial values y(pi/2) = 1 and y'(pi/2) = 0?

OpenStudy (zzr0ck3r):

did you get the last one?

OpenStudy (anonymous):

yea

OpenStudy (zzr0ck3r):

I dont want to do that again if its not helping lol

OpenStudy (zzr0ck3r):

did i mess anyhting up?

OpenStudy (anonymous):

ooh laplace =P

OpenStudy (anonymous):

i didn't get the matrix part, but the rest helped

OpenStudy (anonymous):

but i figure if I can get this last problem done, I should be able to figure all the rest out

OpenStudy (zzr0ck3r):

so are you at the point again where you need to do partial fractions or move some stuff around so your inverse laplaces make since?

OpenStudy (zzr0ck3r):

sense*

OpenStudy (anonymous):

no, I haven't started that yet. honestly, the last few ?'s i posted werent answered withitn ten minutes, so i thought i had some time

OpenStudy (anonymous):

you know how to do this outkast?

OpenStudy (anonymous):

I just learned it yesterday

OpenStudy (anonymous):

hahahaha

OpenStudy (anonymous):

any luck? or any hints on how to start out?

OpenStudy (zzr0ck3r):

aight but give me a sec, it takes way to long for me to type it.

OpenStudy (anonymous):

to start out? hold on let me get my book

OpenStudy (anonymous):

take the laplace of each one

OpenStudy (anonymous):

of the derivatives what do you get

OpenStudy (anonymous):

(s^2-s-2)Y - s - 1 = (-8s/(s^2 + B^2)) - (2B/(s^2+B^2))

OpenStudy (anonymous):

but what would B be?

OpenStudy (anonymous):

\[y''=s^2Y(s)-sy(\pi/2)-y'(\pi/2)=s^2Y(s)-s(1)-0=s^2Y(s)-s\]

OpenStudy (anonymous):

\[y'=sY(s)-y(\pi/2)=sY(s)-1\]

OpenStudy (anonymous):

y=Y(s)

OpenStudy (anonymous):

okay, got the left side correct

OpenStudy (zzr0ck3r):

yeah man your teacher is a d**k the first term expands to 6 terms and then there are 4 more that I have not expanded so its going to be like 12+ terms. sorry man I dont got that much time. I think its just the partial fractions you have a problem with? Dont worry about the initail conditions just call them to variables for now and deal with them at the end. as far as partial fractions I would just google a good tutorail on them, because when you start adding variables to the numerator it gets odd. But i know there are some slight of hand methods that work just fine...google cover up method for partial fraction decompasition.

OpenStudy (anonymous):

the right side is \[-8(\frac{s}{s^2+k^2})+2\frac{k}{s^2+k^2}\]

OpenStudy (anonymous):

it's cool rock, i'm trying to find the easier way here with outkast

OpenStudy (anonymous):

\[(s^2-s-2)Y(s)-s-1=-8(\frac{s}{s^2+k^2}+2(\frac{k}{s^2+k^2})\]

OpenStudy (anonymous):

but i'm confused on what 'k' is

OpenStudy (anonymous):

\[(s^2-s-2)Y(s)=-8(\frac{s}{s^2+k^2}+2(\frac{k}{s^2+k^2})+s+1

OpenStudy (anonymous):

\[(s^2-s-2)Y(s)=-8(\frac{s}{s^2+k^2}+2(\frac{k}{s^2+k^2})+s+1\]

OpenStudy (anonymous):

\[sin(kt)\]

OpenStudy (anonymous):

the form is \[L{sin(kt)}=\frac{k}{s^2+k^2}\]

OpenStudy (anonymous):

but k is a number and not a constant

OpenStudy (anonymous):

\[(s^2-s-2)Y(s)=-8(\frac{s}{s^2+1})+2(\frac{1}{s^2+1})+s+1\]

OpenStudy (anonymous):

divide the whole thing by (s^2-s-2) \[Y(s)=-8(\frac{s}{(s^2+1)(s^2-s-2)})+2(\frac{1}{(s^2+1)(S^2-s-2)})\] \[+\frac{s}{s^2-s-2}+\frac{1}{s^2-s-2}\]

OpenStudy (anonymous):

can it just be the -8?

OpenStudy (anonymous):

one works too

OpenStudy (anonymous):

nevermind, i got it

OpenStudy (anonymous):

not yet, just what to put there

OpenStudy (anonymous):

\[Y(s)=\frac{-8s+2+s(s^2+1)+s^2+1}{(s^2-s-2)(s^2+1}\]

OpenStudy (anonymous):

whaaa? what happened to them all being split up?

OpenStudy (anonymous):

\[Y(s)=\frac{-8s+2+s^3+s+s^2+1}{(s^2-s-2)(s^2+1)}\]

OpenStudy (anonymous):

even if the answer in the back was more like the other formula, with it all split up?

OpenStudy (anonymous):

oooooh

OpenStudy (anonymous):

\[Y(s)=\frac{s^3+s^2-7s+3}{(s^2-s-2)(s^2+1)}\]

OpenStudy (zzr0ck3r):

expanding that is going to be nutty

OpenStudy (anonymous):

split the denominator up into (s-2)(s+1)(s^2+1)

OpenStudy (anonymous):

\[Y(s)=\frac{s^3+s^2-7s+3}{(s-2)(s+1)(s^2+1)}\]

OpenStudy (anonymous):

next use partial fractions \[\s^3+s^2-7s+3=\frac{A}{s-2}+\frac{B}{s+1}+\frac{Cs+D}{s^2+1}\]

OpenStudy (anonymous):

using the cover up method woudl be the wisest way or by expanding

OpenStudy (anonymous):

which method would you like to do

OpenStudy (anonymous):

cover up method?

OpenStudy (anonymous):

ill show you what i mean \[\frac{s^3-s^2-7s+3}{(s-2)(s+1)(s^2+1)}=\frac{A}{s-2}+\frac{B}{s+1}+\frac{Cs+D}{s^2+1}\]

OpenStudy (anonymous):

if you multiply both sides by s-2 you get

OpenStudy (anonymous):

\[\frac{s^3-s^2-7s+3}{(s+1)(s^2+1}=A+\frac{B(s-2}{s+1}+\frac{Cs+D(s-2)}{s^2+1}\]

OpenStudy (anonymous):

letting s=2 wll cancel out the last two terms getting A= something

OpenStudy (anonymous):

s^3 - s^2 - 7s+3 = A(s+1)(s^2 + 1) + B(s-2)(s^2 +1) + ???

OpenStudy (anonymous):

\[\frac{8-4-14+3}{(3)(5)}=A\]

OpenStudy (anonymous):

what is your question?

OpenStudy (anonymous):

if i did it my way, what would I have where C and D are supposed to be?

OpenStudy (anonymous):

what way ar eyou doing?

OpenStudy (anonymous):

are you expanding?

OpenStudy (anonymous):

is that what it's called? then yes

OpenStudy (anonymous):

it looks like you're expanding.. because you have a polynomial of power n, the numerator is always to the power n-1... so since it's a order 2 polynomial where n=2.. your polynomial on the top must be order 1 or 2-1=1

OpenStudy (anonymous):

....sigh, sorry, then ezpanding is the way my teacher is doing it...

OpenStudy (anonymous):

the standard form of a order one is As+B ... but A and B are already used therefore it is Cs+D

OpenStudy (anonymous):

it doesn't matter . these are the two ways of doing it .. If you want to do expansion i can do that way also

OpenStudy (anonymous):

well, if you can just tell me how to expand it so i can then cross multiply everything, I can to the simple math myself

OpenStudy (anonymous):

oh crap, what about the initial values?

OpenStudy (anonymous):

\[\frac{s^3+s^2-7s+3}{(s-2)(s+1)(s^2+1)}=\frac{A}{s-2}+\frac{B}{s+1}+\frac{Cs+D}{s^2+1}\]

OpenStudy (anonymous):

\[s^3+s^2-7s+3=A(s+1)(s^2+1)+B(s-2)(s^2+1)+(Cs+D)(s-2)(s+1)\]

OpenStudy (anonymous):

\[s^3+s^2-7s+3=A(s^3+s+s^2+1)+B(s^3+s-2s^2-2)+(Cs+D)(s^2-s-2)\]

OpenStudy (anonymous):

\[s^3+s^2-7s+3=As^3+As^2+As+A+Bs^3-2Bs^2+Bs-2B+(Cs+D)(s^2-s-2)\]

OpenStudy (anonymous):

wooooh

OpenStudy (anonymous):

ah! can't see the rest

OpenStudy (anonymous):

\[(Cs+D)(s^2-s-2)=Cs^3-Cs^2-2Cs+Ds^2-Ds-2D\]

OpenStudy (anonymous):

and this is the expansion way?

OpenStudy (anonymous):

\[As^3+As^2+As+A+Bs^3-2Bs^2+Bs-2B\] \[+Cs^3-Cs^2-2Cs+Ds^2-Ds-2D=s^3+s^2-7s+3\]

OpenStudy (anonymous):

now you have to group

OpenStudy (anonymous):

\[s^3(A+B+C)+s^2(A-2B-C+D)+s(A+B-2C-D)\] \[+1(A+B-2D)=s^3+s^2-7s+3\]

OpenStudy (anonymous):

\[A+B+C=1\] \[A-2B-C+D=1\] \[A+B-2C-D=-7\] \[A+B-2D=3\]

OpenStudy (anonymous):

now you have to solve all these

OpenStudy (anonymous):

that is why i'd say covermethod is the best way to go here

OpenStudy (anonymous):

@lgbasallote how about this for DE LOL

OpenStudy (anonymous):

so, would you hate me entirely and say no if I asked you to show the cover method?

OpenStudy (anonymous):

....yuuuup! lol

OpenStudy (anonymous):

i Showed you above to find A @iStutts